Aggregator

Autolog save

6 days 18 hours ago
Log Number: 4369922Logbook(s): Tags: Body: saver ran from opsl04

Saved: anthony 05-05-25 16:41:33 {A:5 B:5 C:5 E:1060 A=BS B=BS C=BS D=BS 0R08=0.00 0R08 dump in. A B C dump(s) in.
anthony

Operator In Training SIGN IN

6 days 18 hours ago
Log Number: 4369921Logbook(s): Body:  ACCELERATOR OPERATOR IN TRAINING SIGN IN FORM OPERATOR:    Shriner
DATE:           05/05/2025
SHIFT:             Swing
MENTOR:         Vasilauskis

The operator has read the following:
shriner

ABIL Tag #8669 added to Hall D Tagger Tagger Dump Skid Flow.

6 days 18 hours ago
Log Number: 4369914Logbook(s): Body: The following entry was placed in the Accelerator Bypassed-Interlocks Log (ABIL) Device: Hall D Tagger Tagger Dump Skid Flow Function: FSD for Hall D Aluminum Water Flow Location: FSD Node TG01 Ch. 5 Purpose: Trip Beam if proper flow readback out of limitations/ Flow is good but readback is bad. Elog#4369896 Exp.
webrun

Operator SIGN IN

6 days 18 hours ago
Log Number: 4369913Logbook(s): Body:  ACCELERATOR OPERATOR SIGN IN FORM OPERATOR:    Yoon
DATE:             05/05/2025
SHIFT:               Swing

The operator has read the following:
  • ELOG
  • PD Shift Plan
  • Radiation Surveys for PSS Segments in Controlled Access
  • PSS Log
yoon

Operator SIGN IN

6 days 18 hours ago
Log Number: 4369910Logbook(s): Body:  ACCELERATOR OPERATOR SIGN IN FORM OPERATOR:    Anthony
DATE:             05/05/2025
SHIFT:               Swing

The operator has read the following:
  • ELOG
  • PD Shift Plan
  • Radiation Surveys for PSS Segments in Controlled Access
  • PSS Log
anthony

Vacuum Group Summary 2025-05-05

6 days 19 hours ago
Log Number: 4369905Logbook(s): Body: 

Planned next scroll pump replacements, ion pumps to inspect during Wednesday maintenance day
VTP1L03 scroll pump scheduled for replacement on Wednesday
VIPAD00C
- Is buried deep in Hall D dump, hard to reach
- Is also high radiation area
- Exploring high potting options to fix without involved process of replacement
- More likely need to wait until 2025 SAM, is too involved to get done in a maintenance day

duzik

Hall D Tagger Aluminum Dump Skid

6 days 19 hours ago
Log Number: 4369896Logbook(s): Body: 

As a recap, the skid tripped 3am. The installation crew tried to restart but it would not power up the pump. Facilities verified power was coming to the Sola power supply that powers the relay to turn the pump on. I&C found the Sola to be faulty and replaced it. The pump restarted but the flow switch did not show full flow and eventually reading -0.1 gpm (should be at least 12). Bern restarted the pump a few times and the flow meter reacted the same. Carroll Jones and I verified that the pump is pumping and the discharge pressure is about 75psi.

whitey

CREW CHIEF SIGN IN

6 days 19 hours ago
Log Number: 4369891Logbook(s): Body: 
THOMAS JEFFERSON NATIONAL ACCELERATOR FACILITY
ACCELERATOR OPERATIONS CREW CHIEF SIGN IN FORM

Date : 05-05-25
Time : 15:36:58
Day : Monday
Shift : Swing

Crew Chief : Richardson
Program Deputy: Sereno
Team : X
Arm : Anthony
Designated SSO: Yoon
LERF Operator:
Operators : Anthony, Shriner, Yoon

This stamp certifies that the above Crew Chief has read/reviewed/received the following:
lrichard

Injector work summary

6 days 19 hours ago
Log Number: 4369869Logbook(s): Entry Makers: ywangBody: 

1400-1500 Optimized the injector with beam to ILD:
Figs. 1-4: CW beam from laser A, B, C and D, respectively
Figs. 5-7: B's, C's and D's orbits relative to A's, respectively

ywang

Hall B Halo rates look OK

6 days 19 hours ago
Log Number: 4369868Logbook(s): Entry Makers: mockBody: 

After Injector work, the halo rates in Hall B did not get worse (success!). While optimizing the rest of the injector and removing tail on SLMs.

mock

Testing how injector changes effect Hall B halo rates

6 days 20 hours ago
Log Number: 4369863Logbook(s): Entry Makers: mockBody: 

After Yan made some injector changes. He wanted to check how the halo rates in Hall B would respond. Spoke with the Hall as we told them beam would be off for a longer period of time than they were expecting (sorry!). The Hall agreed to accept some beam to compare halo rates. Starting off at 5 nA and ramping up to 325 nA.

mock

Shift Information

Shift Day
CEBAF Program Physics!!
MCC Crew Chief Blackshaw
MCC Operators Merz, Christian, Janak
Program Deputy Drury
Announcements

Access Status

Recent PSS History

recent pss history graph